# Nonlocal Models in Biology and Life Sciences: Sources, Developments, and Applications

### v1

@@ 2.14 Nonlocal models in mechanobiology and biomechanics
Nonlocal models are essential in mechanobiology and biomechanics for capturing the spatial and temporal aspects of mechanical interactions in biological tissues. These models provide a more comprehensive and realistic representation of mechanical phenomena, offering valuable insights for both research and medical applications. For instance, (i) how mechanical signals are transmitted within cells and tissues, (ii) the influence of mechanical forces on cells and extracellular matrix components, considering the nonlocal effects of stress and strain distribution, (iii) blood vessel mechanics and the effects of blood flow on vessel walls, etc.
Nonlocal beam models have been applied in various engineering and mechanical systems to study the behaviour of slender structures and address challenges that classical local beam theories cannot adequately capture reddy2007 ; park2005 ; liu2006 . They have found applications in various engineering and physics fields, where long-range interactions and size effects influence the behaviour of beams. For instance, bridge-based sensors and actuators have many applications in biological sciences lavrik2004 ; ekinci2005 ; lu2006 . The outcome of these devices depends on their dynamic properties of beamlike elements, which can be improved by studying their micro- or nano-electromechanical system (MEMS or NEMS) structures lu2006 . A set of integro-differential equations can describe the general nonlocal continuum theory, and different kernel functions have been used in the model to incorporate different problems eringen2003 . For example, the equation of motion in nonlocal elasticity theory is given by
@@
where $\rho$ and $f_{l}$ are the mass density and the applied body forces, respectively. For a linear homogeneous elastic solid, the constitutive equations in nonlocal elasticity theory can be written as an integral over the body as lu2006 ; eringen2003 ; eringen1983 :
@@
The operator $L=1-(e_{0}a)^{2}\nabla^{2}$ is a differential operator for a two-dimensional kernel function $\alpha$ satisfying the Green’s function eringen1983 ; lu2007 , and in this case, the constitutive relations simplified to
@@
Researchers have been studying different nonlocal beam models to incorporate different issues, e.g., the nonlocal Timoshenko and Euler beam models lu2006 ; weaver1991 ; zienkiewicz2005 ; wang2005 ; peddieson2003 ; zhang2005 . Flexoelectricity is one of the issues where an electromechanical coupling happens between the polarization and the strain gradient, and these flexoelectric materials play a vital role in making nanoscale sensors, actuators and energy harvesters navickas2006 ; su2020 . Moreover, the nonlocal elasticity theory helps in capturing the long-range interactions and the strain gradient stress of these flexoelectric nanosensors. Taking the nonlocal elasticity theory, the stress and electric displacement tensors as su2020 ; lim2015 ; li2016 :
@@
where $\sigma_{ij}$ and $\sigma_{ijk}$ are the normal and higher-order stress tensors, respectively; $D_{i}$ is the electric displacement vector; $E_{k}$ is the electric field. All the tensors’ expressions are mentioned in su2020 and the references therein. Using these equations, authors have derived a cantilever beam model for the flexoelectric sensor with external loads. Motezaker and his coauthors constructed a sandwich-type nanostructure from a nanoplate covered by two smart layers under the externally applied voltage motezaker2020 . They have considered surface higher-order nonlocal piezoelasticity theory as:
@@
where $\mu_{0}$ and $\mu_{1}$ are the nonlocal parameters; $\sigma_{ij}^{S}$ and $D_{i}^{S}$ surface stress tensor and electric displacement vector, respectively; the remaining terms are mentioned in motezaker2020 and the references therein. Using this higher-order nonlocal model, researchers have shown a decrease in the frequency and bucking load and an increase in the deflection by enhancing the nonlocal parameter motezaker2020 . The opposite happens by applying the negative external voltage.
@@
Nonlocal plate theory is used to evaluate the axisymmetric bending of micro/nanoscale circular plates, accounting for the impact of small-scale influences duan2007 . It is especially pertinent at the micro/nanoscale, where traditional plate theories would not be able to adequately predict the behaviour, as it takes into account the nonlocal effects resulting from the tiny size of the structure. Furthermore, pyroelectricity’s impact on thermally produced vibrations in piezothermoelastic plates is a complex, multidisciplinary issue. To understand the behaviour of such linked systems, researchers frequently employ experimental validation together with numerical models krommer2004 . Variational principles provide a strong and systematic technique for developing the governing equations of motion for piezoelectric, thermopiezoelectric, and hygrothermopiezoelectric materials altay2007 . These concepts are founded on the notion of minimizing or extremizing a certain functional, typically referred to as the total potential energy or the total Lagrangian, with regard to the relevant field variables.
@@
Sometimes, the distances between adjacent metallic structures in a metasurface might get extremely small (near the nanoscale), and in these situations, the classical local response might not be sufficient when taking the spatially independent material characteristics into account. This is due to the quantum nature of the electron gas in metals at such small scales. Such circumstances can be handled via nonlocal effects by adjustments to the resonance frequencies, damping rates, and dispersion relations of the plasmonic modes. This behaviour differs from traditional predictions, especially as the gap size reduces due to the increased prominence of nonlocal influences. Nonlocal effects in plasmonic metasurfaces with nearly contacting surfaces include shifts in resonance frequencies, changes in field enhancement, and changes in near-field distribution yang2020 . These effects are critical for applications such as sensing, imaging, and improved light-matter interactions.
@@
Nonlinear viscoelastic solids refer to materials that exhibit both time-dependent and strain-dependent behaviour under the influence of applied stress or deformation lockett1972 . In this case, the constitutive equations governing the behaviour of nonlinear viscoelastic solids often involve time-dependent and strain-dependent terms csengul2021 . A common representation is through a stress-strain relationship that includes both elastic and viscoelastic components and where the material properties are functions of time and strain. This helps us predict the performance and durability of materials in real-world applications. In addition, studying the thermoelastic behaviour along with viscoelastic solids gives us the experience of stress under the influence of both thermal and mechanical loads loret1990 ; rivera1998 .
@@ 2.15 Peridynamics, fracture and damage problems, multiscale and multiphysics analysis
Peridynamics is a nonlocal theory that has gained popularity in recent years due to its applicability in multiscale and multiphysics research. It is especially effective for modelling materials and structures having discontinuities, such as cracks and fractures, and it provides a framework for considering long-range interactions lejeune2021 . Arguably, the first paper on peridynamics was published in 2000, and the author discussed the models’ discontinuities and long-range forces using a constitutive relation based on pairwise interactions between particles silling2000 . It is being used to capture biological material in fracture, e.g., bone fracture deng2009 ; ghajari2014 . Here, we describe the basic formulation of peridynamics at a spatial point $\mathbf{x}$ and time $t$. Similar to the neighbourhood of a spatial point for nonlocal models, we consider a horizon $\mathcal{H}_{\mathbf{x}}$ of size $\delta_{\mathbf{x}}$ as
@@
Depending on the dimension of the spatial domain, the shape of the horizon $\mathcal{H}_{\mathbf{x}}$ is different, e.g., in one-dimension $\mathcal{H}_{\mathbf{x}}$ is an open interval $(\mathbf{x}-\delta_{\mathbf{x}},\mathbf{x}+\delta_{\mathbf{x}})$; in two dimensions, it is a circular region of radius $\delta_{\mathbf{x}}$ centred at $\mathbf{x}$, etc. In addition, the dual horizon of $\mathbf{x}$ is defined as the union of points whose horizons include $\mathbf{x}$, i.e., $\mathcal{H}_{\mathbf{x}}^{\prime}=\{\mathbf{y}:\mathbf{x}\in\mathcal{H}_{\mathbf{y}}\}$. The equation of motion is formulated as an integral of interaction forces between points on the spatial domain (body) $\Omega$. Following the balance of linear momentum, we obtain lejeune2021 :
${}\rho\ddot{\mathbf{u}}(\mathbf{x},t)=\int_{\mathbf{y}\in\mathcal{H}_{\mathbf{x}}^{\prime}}\mathbf{f}_{\mathbf{x}\mathbf{y}}(\mathbf{\eta},\mathbf{\xi})\mathbf{d}V_{\mathbf{y}}-\int_{\mathbf{y}\in\mathcal{H}_{\mathbf{x}}}\mathbf{f}_{\mathbf{y}\mathbf{x}}(-\mathbf{\eta},-\mathbf{\xi})\mathbf{d}V_{\mathbf{y}}+\mathbf{b}(\mathbf{x},t),$
@@
where $\rho$ is the density, $\ddot{\mathbf{u}}$ is the acceleration of the displacement vector $\mathbf{u}$, $\mathbf{\eta}$ is the relative displacement vector, $\mathbf{\xi}=\mathbf{y}-\mathbf{x}$ is the bond vector between $\mathbf{x}$ and $\mathbf{y}$, $\mathbf{f}_{\mathbf{x}\mathbf{y}}$ is the force density vector per volume acting on particle $\mathbf{x}$ due to particle $\mathbf{y}$, and $\mathbf{b}$ is the body force. Here, the first integration contributes the direct force term acting on the point $\mathbf{x}$, and the second integration contributes the reaction force term. The discrete form of the balance of linear momentum can be obtained by replacing the integrals with summations lejeune2021 . The equation for force density $\mathbf{f}_{\mathbf{x}\mathbf{y}}$ is based on ordinary state-based peridynamics and is a function of collective deformation silling2007 . In computing the force density vectors for different problems, we refer the reader to the peridynamics literature silling2019 ; madenci2014 ; oterkus2015 and the references therein.
@@
Investigating damage problems, particularly in the context of materials with flexoelectricity, adds an additional layer of complexity to the analysis. Flexoelectricity refers to the coupling between strain gradients and electric polarization, and its consideration is crucial for understanding the behaviour of certain materials. The peridynamics formulation for flexoelectricity gives insight into how this nonlocal theory could be developed to incorporate the effects of flexoelectric coupling in materials roy2019 . Many materials have structures on several length scales, and the structural hierarchy can play a significant role in defining bulk material characteristics lakes1993 ; kamath2006 . Furthermore, different stages and challenges have to be faced in understanding and utilizing the flexoelectricity in materials wang2019 . In addition, nonlocal continuum mechanics models are critical for capturing the impacts of long-range interactions and resolving concerns such as material failure, damage, and scale effects. Eringen introduced nonlocal elasticity and gave answers to particular situations such as screw dislocations and surface waves eringen1983 . Nonlocal models account for material inhomogeneities, as well as the effect of gradients in the localization of deformation and fracture singh2020 ; aifantis1992 ; krishnaswamy2020c .
@@ 2.16 Nonlocal boundary and initial conditions and their approximations
Along with the nonlocal model, researchers have used mathematical models with nonlocal boundary and initial conditions bouziani1997 ; merad2015 ; sytnyk2021 , and they play a crucial role in the field of mathematical biology. These conditions can capture the influence of interactions extended beyond the immediate neighbours of many biological processes. For instance, it could represent (i) the migration of species into or out of the study area in ecological models, (ii) the impact of neighbouring cells or tissues at the boundaries of a simulated domain, (iii) the impact of external factors on the initial distribution of a disease when the introduction of the disease happens from outside the modelled region, etc. Here, we first consider a hyperbolic integro-differential equation that models the quasistatic flexure of a thermoelastic rod as merad2015 :
@@
where $f,g,p$, and $q$ are the given functions, $T$ is a positive constant, and $a$ is a suitable defined function that satisfies some conditions merad2015 . This is an example of a nonlocal boundary conditions problem, as the boundary conditions involve all the points in the spatial domain $[0,1]$. Now, we move to the notion of nonlocal initial value problems. We consider a general mathematical model in a bounded domain $\Omega\subset\mathbb{R}^{n}$ ntouyas2006 :
$Lu+c(x,t)u=F(x,t),\leavevmode\nobreak\ \leavevmode\nobreak\ x\in\Omega,0<t<T,$
@@
$u(x,0)+\sum_{k=1}^{N}\beta_{k}(x)u(x,t_{k})=H(x),\leavevmode\nobreak\ \leavevmode\nobreak\ x\in\Omega,t_{k}\in(0,T],k=1,\ldots,N,$
where $L$ is a parabolic operator with continuous and bounded coefficients.

### v2

@@ 2.13 Nonlocal models in mechanobiology and biomechanics
Nonlocal models are essential in mechanobiology and biomechanics for capturing mechanical interactions of the spatial and temporal aspects in biological tissues. These models provide a more comprehensive and realistic representation of mechanical phenomena, offering valuable insights for research and medical applications. For instance, (i) how mechanical signals are transmitted within cells and tissues, (ii) the influence of mechanical forces on cells and extracellular matrix components, considering the nonlocal effects of stress and strain distribution, (iii) blood vessel mechanics and the effects of blood flow on vessel walls, etc.
Nonlocal beam models have been applied in various engineering and mechanical systems to study the behaviour of slender structures and address challenges that classical local beam theories cannot adequately capture [243, 244, 245]. They have found applications in various engineering and physics fields, where long-range interactions and size effects influence the behaviour of beams. For instance, bridge-based sensors and actuators have many applications in biological sciences [60, 61, 62]. The outcome of these devices depends on their dynamic properties of beamlike elements, which can be improved by studying their micro- or nano-electromechanical system (MEMS or NEMS) structures [62]. In addition, comprehensive investigations have been conducted for parabolic and hyperbolic models in deterministic and stochastic setups, taking into account a variety of nonlocal models that arise in the MEMS industry [246, 247, 248, 249, 250]. A set of integro-differential equations can describe the general nonlocal continuum theory, and different kernel functions have been used in the model to incorporate different problems [59, 251]. For example, the equation of motion in nonlocal elasticity theory is given by
@@
where $\rho$ and $f_{l}$ are the mass density and the applied body forces, respectively. For a linear homogeneous elastic solid, the constitutive equations in nonlocal elasticity theory can be written as an integral over the body as [62, 59, 252, 253]:
@@
The operator $L=1-(e_{0}a)^{2}\nabla^{2}$ is a differential operator for a two-dimensional kernel function $\alpha$ satisfying the Green’s function [253, 254], and in this case, the constitutive relations simplified to
@@
Researchers have been studying different nonlocal beam models to incorporate different issues, e.g., the nonlocal Timoshenko and Euler beam models [62, 255, 256, 257, 258, 259]. Flexoelectricity is one of the issues where an electromechanical coupling happens between the polarization and the strain gradient, and these flexoelectric materials play a vital role in making nanoscale sensors, actuators and energy harvesters [260, 261]. Moreover, the nonlocal elasticity theory helps capture the long-range interactions and strain gradient stress of these flexoelectric nanosensors. Taking the nonlocal elasticity theory, the stress and electric displacement tensors as [261, 262, 263]:
@@
where $\sigma_{ij}$ and $\sigma_{ijk}$ are the normal and higher-order stress tensors, respectively; $D_{i}$ is the electric displacement vector; $E_{k}$ is the electric field. All the tensors’ expressions are mentioned in [261] and the references therein. Using these equations, authors have derived a cantilever beam model for the flexoelectric sensor with external loads. Motezaker and his coauthors constructed a sandwich-type nanostructure from a nanoplate covered by two smart layers under the externally applied voltage [264]. They have considered surface higher-order nonlocal piezoelasticity theory as:
@@
where $\mu_{0}$ and $\mu_{1}$ are the nonlocal parameters; $\sigma_{ij}^{S}$ and $D_{i}^{S}$ surface stress tensor and electric displacement vector, respectively; the remaining terms are mentioned in [264] and the references therein. Using this higher-order nonlocal model, researchers have shown a decrease in the frequency and bucking load and an increase in the deflection by enhancing the nonlocal parameter [264]. The opposite happens by applying the negative external voltage.
@@
Nonlocal plate theory is used to evaluate the axisymmetric bending of micro/nanoscale circular plates, accounting for the impact of small-scale influences [265]. It is especially pertinent at the micro/nanoscale, where traditional plate theories would not be able to adequately predict the behaviour, as it takes into account the nonlocal effects resulting from the tiny size of the structure. Furthermore, pyroelectricity’s impact on thermally produced vibrations in piezothermoelastic plates is a complex, multidisciplinary issue. To understand the behaviour of such linked systems, researchers frequently employ experimental validation together with numerical models [266]. Variational principles provide a strong and systematic technique for developing the governing equations of motion for piezoelectric, thermopiezoelectric, and hygrothermopiezoelectric materials [267]. These concepts are founded on minimizing or extremizing a certain functional, typically referred to as the total potential energy or the total Lagrangian, concerning the relevant field variables.
@@
Sometimes, the distances between adjacent metallic structures in a metasurface might get extremely small (near the nanoscale). In these situations, the classical local response might not be sufficient when taking the spatially independent material characteristics into account. This is due to the quantum nature of the electron gas in metals at such small scales. Such circumstances can be handled via nonlocal effects by adjustments to the resonance frequencies, damping rates, and dispersion relations of the plasmonic modes. This behaviour differs from traditional predictions, especially as the gap size reduces due to the increased prominence of nonlocal influences. Nonlocal effects in plasmonic metasurfaces with nearly contacting surfaces include shifts in resonance frequencies, changes in field enhancement, and changes in near-field distribution [268]. These effects are critical for applications such as sensing, imaging, and improved light-matter interactions.
@@
Nonlinear viscoelastic solids refer to materials that exhibit time-dependent and strain-dependent behaviour under the influence of applied stress or deformation [269]. In this case, the constitutive equations governing the behaviour of nonlinear viscoelastic solids often involve time-dependent and strain-dependent terms [270]. A common representation is through a stress-strain relationship that includes both elastic and viscoelastic components and where the material properties are functions of time and strain. This helps us predict the performance and durability of materials in real-world applications. In addition, studying the thermoelastic behaviour along with viscoelastic solids gives us the experience of stress under the influence of both thermal and mechanical loads [271, 272].
@@ 2.14 Peridynamics, fracture and damage problems, multiscale and multiphysics analysis
Peridynamics is a nonlocal theory that has gained popularity in recent years due to its applicability in multiscale and multiphysics research. It is especially effective for modelling materials and structures having discontinuities, such as cracks and fractures, and it provides a framework for considering long-range interactions [68]. The peridynamic theory of solid mechanics substitutes conventional local stress-strain equations with integral formulations in which the force applied to a location is determined by the deformation of the surrounding area. Arguably, the first paper on peridynamics was published in 2000, and the author discussed the models’ discontinuities and long-range forces using a constitutive relation based on pairwise interactions between particles [69]. It is being used to capture biological material in fracture, e.g., bone fracture [66, 273]. Here, we describe the basic formulation of peridynamics at a spatial point $\mathbf{x}$ and time $t$. Similar to the neighbourhood of a spatial point for nonlocal models, we consider a horizon $\mathcal{H}_{\mathbf{x}}$ of size $\delta_{\mathbf{x}}$ as
@@
Depending on the dimension of the spatial domain, the shape of the horizon $\mathcal{H}_{\mathbf{x}}$ is different, e.g., in one-dimension $\mathcal{H}_{\mathbf{x}}$ is an open interval $(\mathbf{x}-\delta_{\mathbf{x}},\mathbf{x}+\delta_{\mathbf{x}})$; in two dimensions, it is a circular region of radius $\delta_{\mathbf{x}}$ centred at $\mathbf{x}$, etc. In addition, the dual horizon of $\mathbf{x}$ is defined as the union of points whose horizons include $\mathbf{x}$, i.e., $\mathcal{H}_{\mathbf{x}}^{\prime}=\{\mathbf{y}:\mathbf{x}\in\mathcal{H}_{\mathbf{y}}\}$. The equation of motion is formulated as an integral of interaction forces between points on the spatial domain (body) $\Omega$. Following the balance of linear momentum, we obtain [68]:
${}\rho\ddot{\mathbf{u}}(\mathbf{x},t)=\int_{\mathbf{y}\in\mathcal{H}_{\mathbf{x}}^{\prime}}\mathbf{f}_{\mathbf{x}\mathbf{y}}(\mathbf{\eta},\mathbf{\xi})\mathbf{d}V_{\mathbf{y}}-\int_{\mathbf{y}\in\mathcal{H}_{\mathbf{x}}}\mathbf{f}_{\mathbf{y}\mathbf{x}}(-\mathbf{\eta},-\mathbf{\xi})\mathbf{d}V_{\mathbf{y}}+\mathbf{b}(\mathbf{x},t),$ (34)
@@
where $\rho$ is the density, $\ddot{\mathbf{u}}$ is the acceleration of the displacement vector $\mathbf{u}$, $\mathbf{\eta}$ is the relative displacement vector, $\mathbf{\xi}=\mathbf{y}-\mathbf{x}$ is the bond vector between $\mathbf{x}$ and $\mathbf{y}$, $\mathbf{f}_{\mathbf{x}\mathbf{y}}$ is the force density vector per volume acting on particle $\mathbf{x}$ due to particle $\mathbf{y}$, and $\mathbf{b}$ is the body force. Here, the first integration contributes the direct force term acting on the point $\mathbf{x}$, and the second integration contributes the reaction force term. The discrete form of the balance of linear momentum can be obtained by replacing the integrals with summations [68]. The equation for force density $\mathbf{f}_{\mathbf{x}\mathbf{y}}$ is based on ordinary state-based peridynamics and is a function of collective deformation [274]. In computing the force density vectors for different problems, we refer the reader to the peridynamics literature [41, 275, 276] and the references therein.
@@
Investigating damage problems, particularly in the context of materials with flexoelectricity, adds an additional layer of complexity to the analysis. Flexoelectricity refers to the coupling between strain gradients and electric polarization, and its consideration is crucial for understanding the behaviour of certain materials. The peridynamics formulation for flexoelectricity gives insight into how this nonlocal theory could be developed to incorporate the effects of flexoelectric coupling in materials [277]. Many materials have structures on several length scales, and the structural hierarchy can play a significant role in defining bulk material characteristics [278, 279]. Furthermore, different stages and challenges have to be faced in understanding and utilizing the flexoelectricity in materials [280]. In addition, nonlocal continuum mechanics models are critical for capturing the impacts of long-range interactions and resolving concerns such as material failure, damage, and scale effects. Eringen introduced nonlocal elasticity and gave answers to particular situations such as screw dislocations and surface waves [253]. Nonlocal models account for material inhomogeneities and the effect of gradients in the localization of deformation and fracture [118, 281, 282].
@@ 2.15 Nonlocal boundary and initial conditions and their approximations
We have discussed so far the problems where the state of a system is specified at the boundaries of the domain or initial conditions for time-dependent problems given as values at a single point in time. An extension of the conventional local boundary and initial conditions used in solving differential equations is the nonlocal boundary and initial conditions [283, 284, 285]. This is especially true for systems in which the boundary conditions are constraints based on the value of the function at the boundary and how it behaves over a larger region, or the initial conditions are based on the system’s past results. These problems play a crucial role in mathematical biology [151]. In addition, they can capture the influence of interactions extended beyond the immediate neighbours of many biological processes. For instance, it could represent (i) the migration of species into or out of the study area in ecological models, (ii) the impact of neighbouring cells or tissues at the boundaries of a simulated domain, (iii) the impact of external factors on the initial distribution of a disease when the introduction of the disease happens from outside the modelled region, etc. Here, we first consider a hyperbolic integro-differential equation that models the quasistatic flexure of a thermoelastic rod as [284]:
@@
where $f,g,p$, and $q$ are the given functions, $T$ is a positive constant, and $a$ is a suitable defined function that satisfies some conditions [284]. This is an example of a nonlocal boundary conditions problem, as the boundary conditions involve all the points in the spatial domain $[0,1]$. Next, we move to the notion of nonlocal initial value problems. We consider a general form of a mathematical model in a bounded domain $\Omega\subset\mathbb{R}^{n}$ as [286]:
$Lu+c(x,t)u=F(x,t),\leavevmode\nobreak\ \leavevmode\nobreak\ x\in\Omega,0<t<T,$ (36)
@@
$u(x,0)+\sum_{k=1}^{N}\beta_{k}(x)u(x,t_{k})=H(x),\leavevmode\nobreak\ \leavevmode\nobreak\ x\in\Omega,$
where $t_{k}\in(0,T],k=1,\ldots,N,$ and $L$ is a parabolic operator with continuous and bounded coefficients. Due to the complexity of analysis and computation, these models require sophisticated mathematical and numerical approaches.